OpenStudy (anonymous):

A local vendor is giving away treats. Two baskets contain 12 bags of chips and 10 candy bars. If you randomly select one treat from each basket, what is the probability that both of the treats are chips?

OpenStudy (anonymous):

a.12/77 b.22/77 c.1/4 d.6/11

OpenStudy (anonymous):

the correct answer is D :)

OpenStudy (anonymous):

do you want me to explain how I got the answer or do you get it?

OpenStudy (anonymous):

can you please explain it

OpenStudy (anonymous):

i really need some help :(

OpenStudy (anonymous):

sure no problem!

OpenStudy (anonymous):

Okay so since there is 2 baskets there will be 24 bags of chips and 20 candy bars so in total there is 44 choices so your denominator is 44.

OpenStudy (anonymous):

the total chances of chips is 24 so you would have 24/44 then you reduce the fraction both 24 and 44 are divisible by 4 so you would get 6/11 do you understand it now?

OpenStudy (anonymous):

yes i get it now thanks for the help :)

OpenStudy (anonymous):

it was wrong it was b.22/77 :(

ganeshie8 (ganeshie8):

\[\large \dfrac{12C2}{22C2} = 2/7 = 22/77\]

OpenStudy (anonymous):

sorry!!! I was sure that it was right I guess it just wasn't meant to be that simple!

OpenStudy (anonymous):

its good thanks for trying to help tho